OpenStudy (artemisxrpg):

Find the inverse of (x^2 - 4)/ 2x^2

OpenStudy (anonymous):

Step one: Replace f(x) or what have you with y. \[f(x)=\frac{x^2-4}{2x^2}\] becomes \[y = \frac{x^2-4}{2x^2}\] Step two: Swap x and y. \[x=\frac{y^2-4}{2y^2}\] Step three: Re-isolate y. I'll start you off by multiplying by 2y^2: \[2y^2x=y^2-4\]

OpenStudy (anonymous):

Meh, I'll just do it now: \[2y^2x-y^2=-4\] We can factor y^2 out of the left hand side, like so: \[y^2(2x-1)=-4\] Now, we can divide by the new quantity on the left. \[y^2 = \frac{-4}{2x-1}\] Finally, to eliminate the exponent on the left hand side, we take the square root of both sides. Remember your plus/minus sign! \[f^{-1}(x)=\pm\sqrt{\frac{-4}{2x-1}}\]

OpenStudy (anonymous):

https://www.desmos.com/calculator/2srcu4irgc Black = original Red = inverse Orange dashed = y=x (the line all inverse functions are reflected over!

OpenStudy (artemisxrpg):

Of course, I had to factor out the y^2. Thanks for the help.